OpenStudy (anonymous):

5 students consisting 2 girls and 3 boys are to be randomly assigned to 5 different chores of which 2 chores are light duty. What is the probability that: a) the girls will be assigned to the light duty chores? b) the girls will not be assigned to the light duty chores?

OpenStudy (anonymous):

Probability= favorable number of events/total nummber events combinations of 5 elements taken 2 by 2 = 5*4/2=10 combinations of 2 elements taken 2 by 2 = 1 a) 1/10 b) 9/10

OpenStudy (anonymous):

you mean 5C2 (from the group of people) and then multiply that with 2C2?

OpenStudy (anonymous):

devide positive combinations by total

OpenStudy (anonymous):

Thanks for replying btw. Wouldnt you go 2C2 (from the girls) instead of the entire group that goes 5C2? would the total be 5! (factorial)?

OpenStudy (anonymous):

but the order doesn't matter

OpenStudy (anonymous):

5C2 =5*4/2!

OpenStudy (anonymous):

oh i see and thats for the total probability?

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

so i think the event would be 2C2 for choosing from the group of girls and from that we go 2C2 again when assigning them the tasks so 2C2.2C2 =1 hence 1/10?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

and for part B its just 1-1/10

OpenStudy (anonymous):

awesome thanks! are you good at perms and combs?

OpenStudy (anonymous):

don't remmeber much about it...:)

OpenStudy (anonymous):

oh ok but thanks for help!